Suppose that you have $50,000 in cash today. You want to find a bank account that offers an interest rate that will allow you to make withdrawals to pay your monthly expenses of $1,000, beginning one month from today, for 5 years before exhausting the account. Assume that any savings account we consider pays interest monthly (i.e., monthly compounding). What APR must the account offer in order for you to achieve your goal? C) 7.42% 47.33% 10.44% 5.00% 8.37%
Answer #1

7.42%

Using financial calculator
Input: PMT=-1000, N=5*12 = 60 , PV= 50000

Find I/Y = 0.618%

APR= 0.618%*12 = 7.42%